A 500cm3 of solution is prepared by dissolving 0.25g of CaCo3 in excess HCl.Ca2+ ions concentration in the solution in ppm?


1
Expert's answer
2022-05-31T07:25:03-0400

CaCO3 + 2 HCl = CaCl2 + CO2 + H2O

n (CaCO3) =0,25g / 100 g/mol = 0,0025 mol;

V = 500sm3 = 500 ml = 0,5 litr.

1mol CaCO3 => 1mol Ca2+

0,0025mol => X

X = 0,0025 mol Ca2+

CM = n / V = 0,0025 mol / 0,5 l = 0,005 mol/l

ANSWER: 0,005 mol/l Ca2+